OpenStudy (anonymous):

Find an equation of each line given one point and the slope (-10,-5) ;slope-5/3

OpenStudy (anonymous):

To solve you could just plug your x and y numbers in to the y=mx+b equation to find your y-intercept.. and then you would have your other equation:) Hope that helped...:)

OpenStudy (anonymous):

Thanks c:

OpenStudy (tkhunny):

"plug" doesn't actually mean anything. You have a Point and a Slope. You should have the Point-Slope Form, usually this: \((y-y_{0})=m(x-x_{0})\) where \((x_{0},y_{0})\) is the given point and \(m\) is the given slope. Make the required substitutions and you will be done.

OpenStudy (anonymous):

Thanks :)